Calculate the impulse necessary to stop a 1500 kg car travelling at 90 km per hour

Grade:11

1 Answers

Sandipkumar
6 years ago
Here m=1500Kg u=90×5/18 mper sec=25v=0Impulse J =change in momentum=m(v-u)=1500×*(25-0) = 37500Kgm/s